HUBBLE SPACE TELESCOPE DAILY REPORT #998 PERIOD COVERED: 0800 EST Daily Status Report as of 294/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 4904 (The Nuclei of Nearby S0 and Spiral Galaxies - IIIA Snapshot Image) The Planetary Camera was used to make one observations of the low-red-shift galaxy NGC-1527 to study its surface brightness profile, identify sub-arc-second scale condensations, find point sources, and look for gasious disk and dust lanes. The observation was executed without problems. The data was received and displayed, and it was reported to look fine. 1.2 Completed PC 4308 (Snapshot Survey for Companions to Planetary Nebula - Cycle 3 High) The Planetary Camera was used to make an observation of the central region of planetary nebula NGC-7009. It is possible that the central star in a double star. The observations was received and displayed in the OSS, the data was reported to look very good. 1.3 Completed HSP/UV2 5070 (Optical and Ultraviolet Observations of Radio Galaxies - Cycle 2 High) The High Speed Photometer was used to make observations of a pulsar which is located in the Large Magellanic Cloud. The UV2 detectors were used, and onboard target acquisition were used on a offset star. The Sample time was 0.0003 second, and the total observation time was about one hour. The OSS monitored the observations, and assessed the data quality. The target acquisition images showed no clear evidence of the star. It was not certain that the target was actually acquired. The science observations were reported to be weak. 1.4 Completed WFC 3589 (Parallel High Resolution Imaging of Defuse Objects in the Magellanic Clouds) The Wide Field Planetary Camera was used to make observations of the nebulous object LMC-N158-P which is located in the Large Magellanic Cloud. This observations was take in parallel with the prime science observation. Am emission line filter and nearby continuum filter were used. Two image were received and displayed. They both were reported to look okay. 1.5 Completed Two sets of WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. Three observations were received and displayed. All three were reported to look normal. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.7 Completed PC 4903 (The Nuclei of Nearby S0 and Spiral Galaxies -II A PC Snapshot Imaging) The Planetary Camera was used to make snapshot images of the low-redshift galaxy IC1919 to study its surface brightness, identify sub-arcsecond condensations, look for point sources, and look for a stellar or gaseous disk and dust lanes. The observation was received and displayed in the OSS. The observations was reported to look okay. 1.8 Completed PC 4929 (WF/PC Cycle 2 - Improved Dark Calibration) The Planetary Camera was used to make one dark calibration image. The exposures lasted 1200 seconds each, and the camera was pre-flashed before the observation. The observation was reported to be a normal exposures. 1.9 Completed FOS/BL 5255 (Multiwavelength Observations of the Flares and Oscillations of AE AQR) The Faint Object Spectrograph (Blue Side) was used to make ultraviolet observations of the cataclysmic variable star AE Aquarii. The binary pair has an orbital period of 9.88 hours. The onboard target acquisition was monitored in real-time. Inspection of the science and engineering data show that the target acquisition was successful. No problems were reported. 1.10 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Significant HSTARs: None 2.4 Operations Note: The SMGT#9 RGA test is being conducted today, Thursday Oct. 21, 1993 from 8am until 8pm. No problems have been reported. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) 3.2 Complete PC 4308 (Snapshot Survey for Companions to Planetary Nebula Nuclei - Cycle 3 High) 3.3 Complete PC 4904 (The Nuclei of Nearby S0 and Spiral Galaxies - III A PC Snapshot Images) 3.4 Complete FOS/BL 4260 (Extreme Stellar Winds and Post Main Sequence Evolution) 3.5 Complete WFC 3589 (Parallel High Resolution Imaging of Defuse Objects in the Magellanic Clouds) 3.6 Complete HSP/UV1/POL 4209 (HSP Detector Sensitivity Test) 3.7 Complete HSP/UV1 4945 (Periodic Variations in DQ Herculis Stars - Cycle 2 High) 3.8 Complete GHRS/2 5040 (GHRS Detector Calibrations - Cycle 3) 3.9 Complete HSP/UV2 4801 (Quasi-Periodic Oscillations in AM Herculis Binaries) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-